Exercise 18-10 Suppose selected comparative statement data for the giant bookseller Barnes & Noble are presented here. All balance sheet data are as of the end of the fiscal year (in millions) 2020 2019 Net sales Cost of goods sold Net income Accounts receivable Inventory Total assets Total common stockholders equity $5,600 3,637 147 109 1,239 3,170 980 $4,800 3,888 1,314 2,830 994 Compute the following ratios for 2020. (Round asset turnover to 2 decimal places, e.g. 1.67. Round all other answers to 1 decimal place, e.g. 1.6 or 1.6%.) (a) (b) (c) (d) (e) Profit margin Asset turnover times Return on assets Return on common stockholders equity Gross profit rate
Answer #1

a) Profit margin = Net income/Net Sales = 77/4800 = 1.6%

b) Asset turnover = Net Sales/Average assets = 4800/3000 = 1.6 Times

c) Return on assets = 77/3000 = 2.6%

d) Return on Common Stockholder's equity = 77/987 = 7.8%

e) Gross profit rate = (4800-3888)/4800 = 19%